# Exploring the Effect of Explanation Content and Format on User Comprehension and Trust

### v1

@@
Abstract.
In recent years, various methods have been introduced for explaining the outputs of “black-box” AI models. However, it is not well understood whether users actually comprehend and trust these explanations. In this paper, we focus on explanations for a regression tool for assessing cancer risk and examine the effect of the explanations’ content and format on the user-centric metrics of comprehension and trust. Regarding content, we experiment with two explanation methods: the popular SHAP, based on game-theoretic notions and thus potentially complex for everyday users to comprehend, and occlusion-1, based on feature occlusion which may be more comprehensible. Regarding format, we present SHAP explanations as charts (SC), as is conventional, and occlusion-1 explanations as charts (OC) as well as text (OT), to which their simpler nature also lends itself. The experiments amount to user studies questioning participants, with two different levels of expertise (the general population and those with some medical training), on their subjective and objective comprehension of and trust in explanations for the outputs of the regression tool. In both studies we found a clear preference in terms of subjective comprehension and trust for occlusion-1 over SHAP explanations in general, when comparing based on content. However, direct comparisons of explanations when controlling for format only revealed evidence for OT over SC explanations in most cases, suggesting that the dominance of occlusion-1 over SHAP explanations may be driven by a preference for text over charts as explanations. Finally, we found no evidence of a difference between the explanation types in terms of objective comprehension. Thus overall, the choice of the content and format of explanations needs careful attention, since in some contexts an explanation’s format, rather than its content, may play the critical role in improving user experience.
Explainable AI, Feature Attribution, User Studies, Trust
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Explainable AI (XAI) has risen to prominence in recent years as a means of giving users the explanations they need to be able to trust the predictions of “black-box” AI models. This trustworthiness of AI models is especially pertinent in high risk settings such as medicine, when patients’ lives may depend on these predictions. The ways in which explainability may facilitate trust in users has been studied extensively (Jacovi et al., 2021; Schoeffer et al., 2022; Ferrario and Loi, 2022; Panigutti et al., 2022), with the general consensus leaning towards it playing an important role in this critical bond between humans and machines. However, despite this growing consensus, what remains to be seen is which forms of XAI, and which requisite characteristics, are best suited to facilitating trust in users in each individual application of AI models. One of the most prominent types of XAI method is feature attribution (FA, see (Bodria et al., 2023) for an overview). FA explanations assign each feature in a given input a positive or negative attribution value denoting its importance towards the AI model’s output. However, blindly applying one FA method or another in a particular situation is not a scientifically rigorous approach to accounting for the human factors involved in enabling trustworthy AI.
The evaluation of XAI methods, particularly those concerning (the often neglected (Keane et al., 2021)) user studies with humans, have thus been of critical importance to ensuring the methods are serving humans’ needs. Various metrics have been studied as dimensions along which to make these evaluations, including those which are machine-centric, e.g. robustness (Jiang et al., 2023), i.e. how sensitive an explanation is to changes in the scenario. However, it is human-centric metrics which present direct opportunities for measuring how humans feel about XAI, such as (perceived) comprehensibility (Schmude et al., 2023), how well users (think they) understand explanations, and trust (Jacovi et al., 2021), i.e. how much users trust explanations. We argue that each individual application setting requires its own analysis along these metrics and optimisation with regards to its unique explanatory requirements, as is called for in (Hamon et al., 2021). Further, very few studies, with some notable exceptions, e.g. (Szymanski et al., 2021; Bertrand et al., 2023; Schmude et al., 2023), have considered the effect of the format, rather than the content, of explanations on these metrics.
@@ 1. Introduction
Concurrently, the application of AI models to the medical domain has been swift and seemingly ubiquitous (see (Tjoa and Guan, 2021) for a recent overview), with suitable areas ranging from image classification for brain tumour detection (Legastelois et al., 2023) to clinical decision support (Domínguez et al., [n.d.]). The huge potential for tangible benefits of AI models to countless human lives means the search for techniques to ensure that the AI models are trustworthy is particularly urgent, and thus XAI methods are in increasing demand here. One recent study (Kostopoulou et al., 2022) investigates whether predictions from a cancer risk prediction model, namely the QCancer algorithm (Hippisley-Cox and Coupland, 2015), have a beneficial effect in supporting the referral decisions of UK-based General Practitioners (GPs). Kostopoulou et al. (2022) found that the inclination of the GPs on whether to refer the patient changed 26% of the time, while their decisions switched entirely only 3% of the time. However, despite the reluctance of GPs to adopt advice from the algorithm, the changed decisions were more consistent with established clinical guidelines, highlighting the potential of this approach to decision support. In a follow-up study, Pálfi et al. (2023) attempted to improve the uptake of the algorithm by providing a naive, chart-based FA explanation that visualised the relative weights (based on the regression coefficients of the model) of the features for each prediction. However, they found no additional benefit of the explanations over providing the prediction without the explanation. They speculated that this failure to enhance the adoption of and trust in the algorithm by providing an XAI method could be related to the difficulty in understanding the explanations. This context thus seems to be relatively fertile ground to examine the potential of established XAI techniques. However, given that the explanatory requirements of AI models within healthcare have been shown to be different from those in general domains (Legastelois et al., 2023), user studies which are tailored to individual contexts are needed, as stated in (Colin et al., 2022).
@@
In this paper, we set out to address some of the unanswered questions outlined above, in particular those highlighted by Pálfi et al. (2023). Namely, we aimed to shed light on differing degrees of (perceived) comprehensibility and trust affected in users when they receive explanations of different types and formats. We do this by deploying different types of XAI methods, namely the FA explanations SHAP (Lundberg and Lee, 2017) and occlusion-1 (Ancona et al., 2018), for the QCancer algorithm, in the realm of prediction algorithms for diagnosis in healthcare. We then undertook two user studies, each with a different group of participants, one sampled from the general population and one from students in medical training, to examine the effect of different users’ explanation goals (Delaney et al., 2023), which has been shown to be an important factor in the evaluation of explanations (Kim et al., 2023). The participants were asked to answer questions in a survey to assess their perceived comprehensibility of and trust in explanations for the outputs of the QCancer algorithm on fictitious patients described by vignettes from (Pálfi et al., 2023). Moreover, we created a novel objective measure of comprehension to control for any potential reporting biases. Differently from existing studies, we also control for the format of explanations (chart or text) to examine their effect on this paradigm. It should be noted, however, that we do not examine text-based SHAP explanations since describing their game-theoretic notions in full would be incomprehensible to most users, and any approximation would be unrepresentative and unfair (we expand on the reasoning for this decision in §3.2). As such, our chosen research questions (RQs) are as follows, all considered within this medical context:
@@
Our expectations were that the null hypothesis would be rejected in all cases (i.e. the RQs would be answered affirmatively), given the additional complexity in SHAP explanations. We were surprised to discover instead that the dominance of occlusion-1 over SHAP explanations along these metrics is unclear and, where present, was possibly due to the use of text format.
The paper is structured as follows. We first overview the literature which is most relevant to our work (§2), then detail the methodology of our studies (§3) and our findings therefrom (§4). We then discuss the results and state the limitations of the studies, before concluding and looking ahead to future work5).
2. Related Work
Of the multitude of works undertaking user studies comparing different types of XAI, one which most closely aligns with ours is that of Bertrand et al. (2023). They deliver FA explanations in different formats to examine textual explanations vs. graphical explanations. They find that the former increases trust relative to the latter, but they posit that this could be due to the explanations’ simplicity. Somewhat in contradiction with other work (Rago et al., 2021), they found that explanations did not improve users’ understanding of models’ recommendations.
Other works made similar conclusions regarding the format of explanations, e.g. Szymanski et al. (2021) assess visual and textual explanations for the predictions of AI models in user comprehensibility, controlling for the expertise of users, as we do in our studies. The authors found that lay users preferred the visual explanations but could easily misinterpret them compared to textual explanations.
Another closely related study (Schmude et al., 2023) evaluates the effect of textual, dialogical and interactive explanations on user understanding. They found that the textual explanations presented greater limits to understanding for and lower preferences from users.
Moving away from the format of explanations, Cheng et al. (2019) compared white- vs. black-box explanations. The authors assessed users’ (perceived) understanding and trust, as we do, validating our decision to differentiate between self-reported and objective comprehensibility. Cheng et al. (2019) found that, among other results related to interactivity, white-box explanations did not increase understanding relative to those which are black-box, validating our decision to use the latter. Surprisingly, however, they also found that education or technical literacy had no effect on these parameters. However, Schoeffer et al. (2022) found that more information gives greater perceptions of information fairness and trustworthiness and that AI literacy also increased both perceptions.
The transparency of explanations is also studied by Bove et al. (2022), who investigate the effect on AI models’ explanations of contextualisation, i.e. giving more transparency by showing what the system does and how it does it, and exploration, i.e. interacting with the explanation/model or navigating different explanations. They assess end users’ understanding of and satisfaction with explanations, failing and succeeding, respectively, to reject the null hypotheses.
Hadash et al. (2022) identify a relevant issue with FA explanations concerning the negative framing of attributions, i.e. the case where a feature with a negative attribution contributes towards a positive decision or vice versa. A user study shows that the positive framing of feature attributions gives significant benefits in terms of user comprehensibility. In this paper we keep the FA explanations in their original framing but plan to investigate positive framing in the future.
Wang and Yin (2023) find that changes to explanations (i.e. lacking robustness) may reduce users’ trust in and satisfaction with the AI model but exclusively in the case where participants have domain knowledge. This demonstrates the importance of separating users by their expertise, as we have done in this paper. Users’ contexts, in the form of their personal relevance, AI expertise and level of stake, were also found to be important factors in designing XAI to inspire trust in the user study of (Yuan et al., 2023). This is backed up by other claims in the literature, e.g. that explanations must be designed to suit the needs of individuals (Bhatt et al., 2020) and also make sense to them (Byrne, 2023).
In summary, there are a number of orthogonal studies to those we undertake here, with many design choices validating those we take, but none consider the assessment of (perceived) comprehensibility and trust of FA explanations, controlling for the explanation format and the participant expertise.
@@ 2. Related Work
Explainability is often considered within the medical domain as a potential tool to enhance trust in and uptake of prediction models (Tjoa and Guan, 2021; Watson et al., 2019). For instance, Lebovitz et al. (2022) conducted a user study in a hospital where AI tools were deployed by radiologists making predictions concerning breast cancer, lung cancer and bone age: they found that trust in the AI models’ predictions, and thus incorporation into the radiologists’ judgements, were highly dependent on their department, pointing to XAI as a potential solution.
In a user study with healthcare providers, Panigutti et al. (2022) assessed the effect of explanations on trust in AI models operating a Clinical Decision Support System. They showed that users preferred explanations from the AI model, though the user satisfaction with the explanations themselves was low. In contrast, a vignette study with intensive care doctors found no additional benefit of an FA explanation over a prediction only (Nagendran et al., 2023). Moreover, doctors, on average, reported somewhat negative attitudes about the usefulness of the explanations presented during the experiment (mean of 1.97 for non-training doctors on a scale from 1 to 5 with 5 being very useful).
In summary, there are few studies that test the potential of XAI methods to improve medical decisions, and they offer mixed results. Several recent reviews (Antoniadi et al., 2021; Nor et al., 2021; Hauser et al., 2022) highlighted these issues and called for further rigorous, empirical evaluation of the potential of XAI methods in healthcare, giving further motivation for our study.
3. Methodology
In this section, we detail our methodology for the user studies. We first cover the collection of the vignettes presented to participants with the predictions made by the QCancer algorithm (§3.1), before defining explanations for the predictions (§3.2). We then detail the experimental design and procedure (§3.3), explain how we recruited participants from the two different populations (§3.4) and outline our method for analysing the data (§3.5).
3.1. Vignettes
@@
We adapted twenty fictitious vignettes (including one practice vignette) from (Pálfi et al., 2023), describing hypothetical patients presenting at the primary care (i.e. their GP) with symptoms and risk factors suggestive of gastro-oesophageal cancer. Note that we chose to focus on one type of cancer here for simplicity, though the QCancer algorithm and the original vignettes concern multiple types of cancer. Each vignette begins with a table including demographics and some risk factors of the patient, and continues with a textual description of a patient‘s symptoms, other risk factors and the results of some relevant examinations. All of the presented factors were predictive of gastro-oesophageal cancer according to the QCancer algorithm. Then, we provided the estimate of the patient’s risk of gastro-oesophageal cancer calculated by the QCancer algorithm followed by an explanation of this prediction and information on how the explanation should be interpreted. An example of one such vignette, with the output of the QCancer algorithm shown for gastro-oesophageal cancer is given in the following (we will give examples of the explanations in §3.2).
@@
Patient name: Joyce Acorn (female), Age: 65, BMI: 30.64 (Height: 177cm, Weight: 96kg), Smoking: Smoker 20 cigarettes per day, Past medical history: Nil, Family history: Nil.
@@
The variables used in the QCancer algorithm’s prediction of gastro-oesophageal cancer and the values which Example 3.1 results in are shown in Table 1 (top and bottom rows, respectively – the table also shows, in the middle, the baseline patient values for males and females, which will be discussed in §3.2).
@@
Table 1. Values for the QCancer algorithm used for the baseline patients and the patient in Example 3.1, with variables including Smoking Status (SS), Type 2 Diabetes (T2), Chronic Pancreatitis (CP), Loss of Appetite (LA), Unintentional Weight Loss (WL), Difficulty Swallowing (DS), Abdominal Pain (AP), Blood in Vomit (BV), Indigestion (I), Heartburn (H), Anaemia (A) and Change in Bowel Habit (BH).
@@
The vignettes were divided into a low (less than 2%) and high (more than 2%) risk sets based on the estimated risk of gastro-oesophageal cancer for the patient computed by the QCancer algorithm. The low risk set ranged from 0.08% to 1.49%, and in the high risk set they ranged from 2.53% to 53.45%. To control for order effects and to avoid the repetition of vignettes during the experiment, the sets of low and high risk vignettes were further divided into two subsets of vignettes with similar risk profiles, giving low risk sets A and B, and high risk sets A and B (see more details in §3.3).
3.2. Explanations for the QCancer Algorithm
@@
In this section, we define the explanations which we deliver to users. To define these explanations we assume a simple, single-label classification problem comprising a classifier $f:\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}\rightarrow[0,1]$, where $\mathcal{D}_{1},\ldots,\mathcal{D}_{k}$ represent the domains for each of the $k$ features ($k\!\!=\!\!15$ in our setting, see Table 1). Then, for any input $\mathbf{x}\!\!\in\!\!\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}$, $f(\mathbf{x})\!\!\in\!\![0,1]$ is the predicted probability of the class given the current values of the input features. This represents the QCancer algorithm, where the features, e.g. age, are inputs to the algorithm and the output is the predicted risk of gastro-oesophageal cancer. We refer to the $i$th value of any input $\mathbf{x}$ as $x_{i}$, where $i\!\!\in\!\!\{1,\ldots,k\}$. In order to define the FA explanations, we first introduce a simplified version of any given input $\mathbf{x}$ as $\mathbf{x}^{\prime}$, where there exists a mapping function $h_{\mathbf{x}}$ such that $\mathbf{x}\!\!=\!\!h_{\mathbf{x}}(\mathbf{x}^{\prime})$, i.e. returning the original input given the simplified version. We then let, as in (Lundberg and Lee, 2017), $\mathbf{x}^{\prime}\!\!\in\!\!\{0,1\}^{M}$, where $M$ is the number of simplified input features (note that $M\!\!=\!\!k$ throughout this paper). Then, if $x_{i}^{\prime}\!\!=\!\!1$, then $\mathbf{x}\!\!=\!\!h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the current value of the feature, while if $x_{i}^{\prime}\!\!=\!\!0$, then $\mathbf{x}\!\!=\!\!h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the feature’s baseline value, i.e. a pre-defined default value. We set the baseline values (referred to as baseline patients in the explanations) to one of two sets of values representing healthy males and females depending on the patient’s gender, as indicated in Table 1, which were used for the studies.
@@
One of the most prominent of all FA methods is SHAP (Lundberg and Lee, 2017), where each feature of an input to a model is assigned an importance value computed from a game-theoretical approximation of its contribution.
Definition 3.2.
@@
where $\mathbf{z}^{\prime}\subseteq\mathbf{x}^{\prime}$ represents that $\mathbf{z}^{\prime}$ is a vector whose non-zero entries are a subset of the non-zero entries in $\mathbf{x}^{\prime}$, $|\mathbf{z}^{\prime}|$ is the number of non-zero entries in $\mathbf{z}^{\prime}$ and $\mathbf{z}_{-i}^{\prime}$ denotes $\mathbf{z}^{\prime}$ after setting $z_{i}^{\prime}=0$.
@@
We then introduce chart-based SHAP (SC) explanations, which are the explanations computed by SHAP expressed as a percentage in chart format, as exemplified in Figure 1 (bottom). [^1] On acceptance, we will share the link to our explanation visualisation tool. Note that, for all explanations, the displayed features were the six with the highest magnitudes. Note also that all explanations were accompanied by a textual description of the explanation, as shown in Figure 1 (top). In the example explanation in Figure 1, we can see that the top five features increased the risk of gastro-oesophageal cancer, with difficulty swallowing and its value of yes being the most important. The last feature, BMI, with value 30.643, reduced the risk but had a smaller magnitude, and thus importance, as calculated by SHAP.
Figure 1. A SC explanation (bottom) for the vignette given in Example 3.1, and textual description of the SC explanation (top). Red (blue) bars represent features which increased (decreased, respectively) the risk of cancer. The patient’s values are given in bold and the baseline values in grey.
We then selected the occlusion-1 method (Ancona et al., 2018) for its simplicity in hopes of achieving more comprehensibility and trust in users. This method takes the set of baseline feature values (as in the SHAP explanations) and assigns any feature an attribution corresponding to the difference in the prediction if the feature’s value is changed to the baseline value (with all other features’ values remaining unchanged).
Definition 3.3.
@@
Note that, while these explanations do not explicitly mention datapoints, each feature’s attribution in explanations implicitly refers to a modified datapoint with that feature set to the baseline value, with all other features remaining unchanged. Thus, while this differs from counterfactual and semi-factual explanations in the literature (see (Aryal and Keane, 2023; Guidotti, 2022), respectively, for recent overviews [^2] Also that there are works which draw important links between FA and counterfactual explanations (Mothilal et al., 2021; Albini et al., 2022).), which are based on single modified datapoints, it still gives an actionable representation of (six) modifications users could make to the input to achieve a different output (similarly to Albini et al. (2020)). We therefore tentatively predicted that this inherent counterfactual nature of occlusion-1 explanations may make them more amenable to humans, as observed for counterfactual explanations by Miller (2019), which strengthened our prediction of the rejection of the null hypotheses for RQ1-3.

### v2

@@
Abstract
AI-driven tools for healthcare are widely acknowledged as potentially beneficial to health practitioners and patients, e.g. the QCancer regression tool for cancer risk prediction. However, for these tools to be trusted, they need to be supplemented with explanations. We examine how explanations’ content and format affect user comprehension and trust when explaining QCancer’s predictions. Regarding content, we deploy SHAP and Occlusion-1. Regarding format, we present SHAP explanations, conventionally, as charts (SC) and Occlusion-1 explanations as charts (OC) as well as text (OT), to which their simpler nature lends itself. We conduct experiments with two sets of stakeholders: the general public (representing patients) and medical students (representing healthcare practitioners). Our experiments showed higher subjective comprehension and trust for Occlusion-1 over SHAP explanations based on content. However, when controlling for format, only OT outperformed SC, suggesting this trend is driven by preferences for text. Other findings corroborated that explanation format, rather than content, is often the critical factor.
1 Introduction
The deployment of AI-driven tools in the healthcare domain has been widespread (see Tjoa and Guan (2021) for an overview), in areas from image classification for brain tumour detection Legastelois et al. (2023) to clinical decision support Domínguez et al. (2023). However, to foster trust in users, the underlying “black-box” AI models need to be supplemented with explanations with consideration of the types of stakeholders the users represent Langer et al. (2021). Explainable AI (XAI) has risen to prominence in recent years as a means of providing users with the explanations they need to be able to trust the outputs of AI models. The ways in which explainability may facilitate trust in users has been studied extensively Ferrario and Loi (2022); Ma et al. (2023), with the general consensus leaning towards it playing an important role in this bond between humans and machines. However, what remains to be seen is which forms of XAI are best suited in individual applications of AI models Hamon et al. (2021). Further, given that the explanatory requirements within healthcare are different from those in general Legastelois et al. (2023), such tailored deployments are especially important in this domain Corti et al. (2024).
One prominent XAI method is feature attribution (FA, see Bodria et al. (2023) for an overview). FA explanations assign each feature for a given input an attribution value denoting its importance towards the output. However, FA methods alone do not account for the human factors involved in enabling trustworthy AI. To evaluate how well XAI methods perform wrt these factors are human-centric metrics such as (perceived) comprehensibility Schmude et al. (2023), i.e. how well users (think they) understand explanations, and trust Jacovi et al. (2021), i.e. how much users trust explanations. However, only a limited number of studies, e.g. Szymanski et al. (2021); Bertrand et al. (2023); Schmude et al. (2023), have considered the effect of the format, rather than the content, of explanations on these metrics.
@@ 1 Introduction
One recent study Kostopoulou et al. (2022) investigates whether predictions from a cancer risk prediction model, namely the QCancer algorithm Hippisley-Cox and Coupland (2015), have a beneficial effect in supporting the referral decisions of UK-based General Practitioners (GPs). The study found that the inclination of the GPs on whether to refer the patient changed 26% of the time, while their decisions switched entirely only 3% of the time. However, despite the reluctance of GPs to trust and adopt advice from the algorithm, the changed decisions were more consistent with established clinical guidelines, highlighting the potential of this approach to decision support. In a follow-up study, Pálfi et al. Pálfi et al. (2023) found that providing a naive, chart-based FA explanation that visualised the relative weights (based on the regression coefficients of the model) of the features for each prediction did not improve the trust in and uptake of the algorithm, speculating that this could be related to the GPs’ difficulty in comprehending the explanations.
@@
In this paper, we set out to address some of the unanswered questions highlighted by Pálfi et al. Pálfi et al. (2023), broadening from healthcare practitioners to include potential patients also. Namely, we undertook a study involving different stakeholders, aiming to shed light on differing degrees of (perceived) comprehensibility and trust affected in users when they receive explanations of different contents and formats. We do this, in collaboration with healthcare practitioners, by deploying different, well-known FA explanation methods, SHAP Lundberg and Lee (2017) and Occlusion-1 Ancona et al. (2018), for the QCancer algorithm. We undertook two user studies, each with a different group of participants, one sampled from the general population and one from students in medical training, to examine the effect of different stakeholders’ expertise and explanation goals Delaney et al. (2023), which are important factors in the evaluation of explanations Kim et al. (2023); Ehsan et al. (2024). The participants were asked to take part in a survey to assess their (perceived) comprehensibility of and trust in explanations for the outputs of the QCancer algorithm on fictitious patients described by vignettes from Pálfi et al. (2023). Moreover, we created a novel objective measure of comprehension (differentiating from perceived comprehension) to control for any potential reporting biases. Differently from existing studies, we control for the format of explanations (chart or text) to examine their effect on this paradigm. We thus aim to explore the effect of explanation content and format on users’ comprehension and trust via the following research questions (RQs), all considered within the healthcare context:
@@
Our expectations were for the null hypothesis to be rejected (i.e. the RQs would be answered affirmatively), given the complexity in SHAP explanations. We were surprised to discover instead that the dominance of the simpler Occlusion-1 explanations is unclear and, where present, may be due to the use of the textual format. These findings provide guidance for the future design of AI-based decision support in healthcare.
2 Related Work
User Studies Comparing Types of XAI
Some works, e.g. Bertrand et al. (2023), also assess FA explanations in text and chart formats, finding that the former increase trust relative to the latter, possibly due to their simplicity. Others have reached similar conclusions, e.g. Szymanski et al. Szymanski et al. (2021) assess visual and textual explanations wrt comprehensibility, controlling for user expertise, as we do, but finding that lay users prefer visual explanations while often misinterpreting them compared to textual ones. Finally, Schmude et al. Schmude et al. (2023) evaluate user comprehension of textual, dialogical and interactive explanations, finding that the former result in worse comprehension and lower preferences from users.
Many works show the importance of expertise with regard to user comprehension and trust, e.g. Wang and Yin (2023), and call for explanations to be designed to suit the needs of individuals Bhatt et al. (2020) and make sense to them Byrne (2023). Other works validate our design choices, e.g. our decision to differentiate between self-reported and objective comprehensibility Cheng et al. (2019). Our approach is a step towards the effective design of multi-modal explanations, e.g. as in Szymanski et al. (2023) in the healthcare domain.
None of these works assess (perceived) comprehensibility and trust of FA explanations, controlling for the format and the participant expertise, as we do.
User Studies of XAI in Healthcare
@@ 2 Related Work
XAI is often considered within healthcare as a tool to enhance trust in and uptake of prediction models Tjoa and Guan (2021). Lebovitz et al. Lebovitz et al. (2022) conducted a user study in which AI was deployed by radiologists to predict breast cancer, lung cancer and bone age: they found that trust in the AI models’ predictions, and thus incorporation into the radiologists’ judgements, were highly dependent on their department, pointing to customised XAI as a potential solution. Moreover, a recent vignette-based user study found promising results Naiseh et al. (2023) in XAI’s deployment in healthcare, finding the AI to be more comprehensible when they received an explanation compared to when they did not. Also, Lee et al. Lee et al. (2024) showed that interactive explanations increased healthcare professionals’ trust in AI, along with their decision accuracy.
Some studies obtained less promising results. For instance, in a user study with healthcare providers, Panigutti et al. Panigutti et al. (2022) assessed the effect of explanations on trust in AI operating a clinical decision support system. They showed that users preferred explanations from the AI model, though the user satisfaction with the explanations themselves was low. Wysocki et al. Wysocki et al. (2023) found that when receiving explanations for COVID-19 risk predictions, users found FA charts easy to interpret but there was no evidence that providing them increased trust in and satisfaction with the AI, compared to when receiving predictions without explanations. A vignette-based study with intensive care doctors found no benefit of an FA explanation over a prediction only option for prescription decisions Nagendran et al. (2023). Finally, a mixed-methods, vignette-based study found that chart-based FA and text-based counterfactual explanations failed to meet the expectations of physicians in reducing uncertainty around heart disease diagnosis Lesley and Hernández (2024).
In summary, few works test XAI’s benefits for healthcare with user studies, with mixed results, which is a known deficiency in the field Hauser et al. (2022); Sadeghi et al. (2024).
3 Methodology
Vignettes
@@ 3 Methodology
We adapted twenty fictitious vignettes (including one practice vignette) from Pálfi et al. (2023), describing hypothetical patients presenting at the primary care (i.e. their GP) with symptoms and risk factors suggestive of gastro-oesophageal cancer. We chose to focus on one type of cancer here for simplicity, though the QCancer algorithm and the original vignettes concern multiple types. Each vignette begins with a table of patient demographics and some risk factors, and continues with a textual description of the patient’s symptoms, other risk factors and the results of some relevant examinations. All of the presented factors were predictive of gastro-oesophageal cancer according to the QCancer algorithm. Then, we provided the estimate of the patient’s risk of gastro-oesophageal cancer calculated by the QCancer algorithm with an explanation of this prediction and information on how the explanation should be interpreted. An example of a vignette, with the output of the QCancer algorithm shown for gastro-oesophageal cancer, is given in the following.
@@
Patient name: Joyce Acorn (female), Age: 65, BMI: 30.64 (177cm, 96kg), Smoking: 20 cigarettes per day, Past medical history: Nil, Family history: Nil.
@@
The variables used in the QCancer algorithm’s prediction of gastro-oesophageal cancer and the values in Example 1 are shown in Table 1, along with the the baseline patient values for males and females, which will be discussed next.
@@
Table 1: Values for the QCancer algorithm used for the baseline patients and the patient in Example 1, with variables including Smoking Status (SS), Type 2 Diabetes (T2), Chronic Pancreatitis (CP), Loss of Appetite (LA), Unintentional Weight Loss (WL), Difficulty Swallowing (DS), Abdominal Pain (AP), Blood in Vomit (BV), Indigestion (I), Heartburn (H), Anaemia (A) and Change in Bowel Habit (BH).
@@
The vignettes were divided into a low and high risk sets based on the estimated risk of gastro-oesophageal cancer for the patient computed by the QCancer algorithm and rounded to the closest integer. We based the threshold for high-risk cases on the 3% referral threshold for suspected cancer of the NICE guidelines for UK healthcare professionals [^2] https://www.nice.org.uk/guidance/ng12/evidence/full-guideline-pdf-2676000277. To control for order effects and to avoid the repetition of vignettes during the experiment, the sets of low and high risk vignettes were further divided into two subsets of vignettes with similar risk profiles, giving low risk sets A and B, and high risk sets A and B (more details are given later in this section).
Explanations for the QCancer Algorithm
@@
To define the explanations that we deliver to users we assume a simple, single-label classification problem comprising a classifier $f:\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}\rightarrow[0,1]$, where $\mathcal{D}_{1},\ldots,\mathcal{D}_{k}$ represent the domains for each of the $k$ features ($k=15$ in our setting, see Table 1). Then, for any input $\mathbf{x}\in\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}$, $f(\mathbf{x})\in[0,1]$ is the predicted probability of the class given the current values of the input features. This represents the QCancer algorithm, where the features, e.g. age, are inputs to the algorithm and the output is the predicted risk of gastro-oesophageal cancer. We refer to the $i$th value of any input $\mathbf{x}$ as $x_{i}$, where $i\in\{1,\ldots,k\}$. In order to define the FA explanations, we first introduce a simplified version of any given input $\mathbf{x}$ as $\mathbf{x}^{\prime}$, where there exists a mapping function $h_{\mathbf{x}}$ such that $\mathbf{x}=h_{\mathbf{x}}(\mathbf{x}^{\prime})$, i.e. returning the original input given the simplified version. We then let, as in Lundberg and Lee (2017), $\mathbf{x}^{\prime}\in\{0,1\}^{M}$, where $M$ is the number of simplified input features (note that $M=k$ throughout this paper). Then, if $x_{i}^{\prime}=1$, then $\mathbf{x}=h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the current value of the feature, while if $x_{i}^{\prime}=0$, then $\mathbf{x}=h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the feature’s baseline value, i.e. a pre-defined default value. We set the baseline values (referred to as baseline patients in the explanations) to one of two sets of values representing healthy males and females depending on the patient’s gender, as indicated in Table 1, which were used for the studies.
@@
One of the most prominent of all FA methods is SHAP Lundberg and Lee (2017), where each feature of an input to a model is assigned an importance value computed from a game-theoretical approximation of its contribution.
Definition 1.
@@
where $\mathbf{z}^{\prime}\subseteq\mathbf{x}^{\prime}$ means that $\mathbf{z}^{\prime}$ is a vector whose non-zero entries are a subset of those in $\mathbf{x}^{\prime}$, $|\mathbf{z}^{\prime}|$ is the number of non-zero entries in $\mathbf{z}^{\prime}$ and $\mathbf{z}_{-i}^{\prime}$ denotes $\mathbf{z}^{\prime}$ after setting $z_{i}^{\prime}=0$.
@@
We then introduce chart-based SHAP (SC) explanations, which are the explanations computed by SHAP expressed as a percentage in chart format, as exemplified in Figure 1 (bottom). [^3] We will share the link to our visualisation tool on acceptance. For all explanations, the displayed features were the six with the highest magnitudes. Note that all explanations were accompanied by a textual description of the explanation, as shown in Figure 1 (top). In the example explanation in Figure 1, we can see that the top five features increased the risk of gastro-oesophageal cancer, with difficulty swallowing and its value of yes being the most important. The last feature, BMI, with value 30.643, reduced the risk but had a smaller magnitude, and thus importance, as calculated by SHAP.
Figure 1: SC explanation (bottom) for the vignette given in Example 1, and textual description thereof (top). Red (blue) bars represent features which increased (decreased, respectively) the risk of cancer. The patient’s values are given in bold and the baseline values in grey.
We then selected the Occlusion-1 method Ancona et al. (2018) for its simplicity in hopes of achieving more comprehensibility and trust in users. This method takes the set of baseline feature values (as in SHAP explanations) and assigns any feature an attribution corresponding to the difference in the prediction if the feature’s value is changed to the baseline value (with all other features’ values remaining unchanged).
Definition 2.
@@
Note that, while these explanations do not explicitly mention datapoints, each feature’s attribution in explanations implicitly refers to a modified datapoint with that feature set to the baseline value, with all other features remaining unchanged. Thus, while this differs from counterfactual and semi-factual explanations in the literature (see Guidotti (2022); Aryal and Keane (2023), respectively, for recent overviews), which are based on single modified datapoints, it still gives an actionable representation of (six) modifications users could make to the input to achieve a different output (as in Albini et al. Albini et al. (2020)). We thus tentatively predicted that this inherent counterfactual nature of Occlusion-1 explanations may give more amenability to humans, as observed for counterfactual explanations by Miller Miller (2019), which strengthened our prediction of the rejection of the null hypotheses for RQ1-3.